OpenStudy (anonymous):

For What Value(s) of x does the expression below equal 0? (x-2)(x+1)/ x(x+3)

OpenStudy (zarkon):

set the numerator equal to zero and solve

OpenStudy (anonymous):

How ?

OpenStudy (zarkon):

solve \[(x-2)(x+1)=0\]

OpenStudy (anonymous):

what about the denominator

OpenStudy (zarkon):

if you have \[\frac{f(x)}{g(x)}=0\] if you multiply by g(x) to both sides you get \[f(x)=0\] so the zeros are determined by the numerator (provided that g(x) is not zero at the same time) also, if you have a product of functions equal to zero set each equal to zero and solve so \[(x−2)(x+1)=0\Rightarrow (x-2)=0\text{ or }(x+1)=0\]

OpenStudy (anonymous):

so thats it or is there more to it?